OpenStudy (anonymous):

f(x)=(5X+1)/(4-3x) 1.what is its domain? 2.does it have an inverse? if it does, what is the domain of the inverse(that would be the range of f)

OpenStudy (anonymous):

i think one of the domain is X not equal to 4/3

OpenStudy (anonymous):

set the denominator = to zero (4-3x)=0 4/3=x so its range is all reals except were x = 4/3

OpenStudy (anonymous):

\[x \in R, x \neq 4/3\]

OpenStudy (anonymous):

does it have an inverse?

OpenStudy (anonymous):

\[f^{-1}(x)=\frac{4x-1}{5+3x}\]

myininaya (myininaya):

(1) the domain is all reals except x=4/3 (2) \[y=\frac{5x+1}{4-3x}\] \[(4-3x)y=4x+1\] \[4y-3xy=4x+1\] \[-3xy-4x=1-4y\] \[x(-3y-4)=1-4y\] \[x=\frac{1-4y}{-3y-4}\] \[f^{-1}(x)=\frac{1-4x}{-3x-4}\] domain of f inverse is all read numbers except when -3x-4=0 (x=-4/3) so the range of f is all real numbers except x=-4/3

OpenStudy (anonymous):

u did a mistake myininaya

OpenStudy (anonymous):

(4-3x)y=5x+1 not 4x+1

myininaya (myininaya):

probably my screen wouldn't stay still the latex was making it go crazy

myininaya (myininaya):

i see it lol you are right i wrote down the correct thingy and then i wrote down the wrong thingy oh well at least it doesn't change the domain and range

OpenStudy (anonymous):

yes but the inverse function is changed

myininaya (myininaya):

oh wait it does change it we hav3e -3x-5 in the bottom not -3x-4

myininaya (myininaya):

-3x-5=0 when x=-5/3 so domain of f inverse is all real numbers but x=-5/3 so the range of f is all real numbers but x=-5/3

OpenStudy (anonymous):

yes..now thats rigth myininaya . i have done it u can check..

myininaya (myininaya):

no its cool its late and i should go to bed

OpenStudy (anonymous):

thank you everyone for your generous help